OpenStudy (anonymous):

Simplify into 1 term: [((sin(x-y))/(cos(x-90)sin(90-y)))-(cos2x + sin2xsinxsecx)] cotx the numbers are in degrees and sorry if I confused things with the parentheses!

OpenStudy (anonymous):

aye is that cos(2x/0 r cos ^2 x

OpenStudy (anonymous):

cos(2x) sry i meant

OpenStudy (lalaly):

\[(\frac{\sin(x-y)}{\cos(x-90)\sin(90-y)}-\cos2x+\sin2xsinxsecx)cotx\]is this it?

OpenStudy (anonymous):

sorry! yest this is and how do you know it is cos(2x)?

OpenStudy (anonymous):

there are parentheses between the nrgative and cos2x

OpenStudy (anonymous):

*negative

OpenStudy (callisto):

Most probably it's wrong :|

OpenStudy (anonymous):

this is what I got but one term only :(

OpenStudy (callisto):

It's one term..... :P

OpenStudy (callisto):

xy = one term.. xy + z = 2 terms

OpenStudy (anonymous):

oh omg I'm an idiot -____- thank you!

OpenStudy (callisto):

welcome ! It might be wrong though.... Sorry for poor hand-writing!!!